Scott earns $6 an hour. He has $15. Using m=15+6h how any hours until he has earned $51?
Thank you for steps!!

Found 2 solutions by addingup, fractalier:
Answer by addingup(3677)

I just answered the same question for someone else, except the character in the other problem was a girl:
-----------------------------------
If she has 15 and makes 6/hr:
6h+15 = 51
6h = 36
h = 6 it will take her 6 work hours to have 51

Answer by fractalier(6550)

Here we set m = 51 and then solve for the number of hours, h. So we write
m = 15 + 6h
51 = 15 + 6h
Now subtract 15 and then divide by 6 and get
36 = 6h
h = 6 hours